# Simultaneous spatial-parametric collocation approximation for parametric PDEs with log-normal random inputs

### v3

@@ 3 Applications to holomorphic functions
We recall a concept of $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions which has been introduced in [26, Definition 4.1]. For $m\in{\mathbb{N}}$ and a positive sequence ${\boldsymbol{\varrho}}=(\varrho_{j})_{j=1}^{m}$, we put
@@
The following key result on weighted $\ell_{2}$-summability of $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions has been proven in [26, Corollary 4.9].
@@
Assume that $\boldsymbol{b}$} is a non-increasing sequence and that $b_{j}^{p-1}\frac{\xi}{4\sqrt{\eta!}}\left\|{{\boldsymbol{b}}}\right\|_{\ell_{p}({\mathbb{N}})}>1$ for all $j\in{\mathbb{N}}$. Then we have
@@
Let $v$ satisfy Assumption 3.1. Let $q_{i}:=p_{i}/(1-p_{i})$ for $i=1,2$, and $\varepsilon$ be any positive number. Let the number $\beta$ be defined as in (2.22). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.29). Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\mathbb{R}}^{\infty}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.26)–(2.28),
@@ 4 Multi-level sparse-grid interpolation algorithms
On the other hand, it holds the following statement which can be proven in a way similar to the proof of [24, Theorem 3.8]. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that
@@ 5 Applications to parametric elliptic PDEs
where $\nu$ is a constant depending on $D$ and $r$. The following result has been proven in [26, Theorem 7.8]
@@
We extend the operators $P_{k}$ (which with an abuse is denoted again by $P_{k}$) to $L_{2}({\mathbb{R}}^{\infty},V;\gamma)$ by the formula
@@
$b_{1,j}:=\left\|{\psi_{j}}\right\|_{L^{\infty}},\quad b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\}$ (5.4)
@@
Let $D\subset{\mathbb{R}}^{2}$ be a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$. Let $\alpha=\frac{r-1}{2}$. Let $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$, and $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$. Let $0<p_{1},p_{2}<1$. For $i=1,2$, let the sequences ${\boldsymbol{b}}_{i}:=(b_{i,j})_{j\in{\mathbb{N}}}$ be defined as in (5.4) with ${\boldsymbol{b}}_{i}\in\ell^{p_{i}}({\mathbb{N}})$. Then there exist $\xi>0$ and $\delta>0$ such that for the parametric solution to (5.1) with inputs (1.3)
@@
Proof. The proof of this lemma is a modification of the proof of [26, Proposition 7.12]. We apply [26, Theorem 4.11] on holomorphy of composite functions with $E=L^{\infty}(D)$ and $X=V$ to prove the claims (1)–(5).
Step 1. We verify the claim (1). By [26, Proposition 7.12] there exist $\xi_{1}>0$ and $\delta_{1}>0$ such that $u({\boldsymbol{y}})$ is $({\boldsymbol{b}}_{1},\xi_{1},\delta_{1},V)$-holomorphic on the open set
$O_{1}=\{a\in L^{\infty}(D;{\mathbb{C}})\,:\,\rho(a)>0\}\subset L^{\infty}(D;{\mathbb{C}})$
@@
Step 2. Concerning the claim (2), by assumption, $b_{1,j}=\left\|{\psi_{j}}\right\|_{L^{\infty}}$ satisfies ${\boldsymbol{b}}_{1}=(b_{1,j})_{j\in{\mathbb{N}}}\in\ell^{p_{1}}({\mathbb{N}})\subseteq\ell^{1}({\mathbb{N}})$, which corresponds to the assumption (iv) of Theorem [26, Theorem 4.11]. It remains to verify the assumptions (i), (ii) and (iii) of [26, Theorem 4.11] for $P_{k}u$, $k\in{\mathbb{N}}_{0}$.
@@
(iii) For all $a$, $b\in O_{1}$ we have by Lemma 5.2 and [26, (4.21)],
$\left\|{P_{k}{\mathcal{U}}(a)-P_{k}{\mathcal{U}}(b)}\right\|_{V}\leq C\left\|{{\mathcal{U}}(a)-{\mathcal{U}}(b)}\right\|_{V}\leq\left\|{f}\right\|_{H^{-1}(D)}\frac{1}{\min\{\rho(a),\rho(b)\}^{2}}\left\|{a-b}\right\|_{L^{\infty}}.$ (5.6)
According to [26, Theorem 4.11] the map
${\mathcal{U}}\mapsto P_{k}{\mathcal{U}}\in L^{2}(U,V;\gamma)$
@@
and verify again the assumptions (i), (ii) and (iii) of [26, Theorem 4.11] with $E=W^{1}_{\infty}(D)\cap{\mathcal{W}}^{s-1}_{\infty}(D)$. First, observe that with
$b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{{\mathcal{W}}^{s-1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}}(D)\big\},$
@@
${\boldsymbol{b}}_{2}=(b_{2,j})_{j\in{\mathbb{N}}}\in\ell^{p_{2}}({\mathbb{N}})\hookrightarrow\ell^{1}({\mathbb{N}})$
which corresponds to the assumption (iv) of Theorem [26, Theorem 4.11].
@@
$\left\|{({\mathcal{U}}-P_{k}{\mathcal{U}})(a)}\right\|_{V}\leq\delta_{k}\frac{(\left\|{a}\right\|_{W^{1}_{\infty}(D)}+\left\|{a}\right\|_{{\mathcal{W}}^{r-1}_{\infty}(D)})^{N_{r}+1}}{\rho(a)^{N_{r}+2}},$
where $\delta_{k}:=K2^{-\alpha k}\left\|{f}\right\|_{{\mathcal{K}}_{\varkappa-1}^{r-2}}$.
(iii) For every $a$, $b\in O_{2}\subseteq O_{1}$, by (5.6) and [26, (4.21)],
@@
$\leq C\left\|{f}\right\|_{H^{-1}(D)}\frac{2}{\min\{\rho(a),\rho(b)\}^{2}}\left\|{a-b}\right\|_{L_{\infty}(D)}.$
We conclude with [26, Theorem 4.11] that there exist $\xi_{4}$ and $C_{4}$ depending on $O_{2}$, $D$ but independent of $k$ such that $u-P_{k}u$ is $({\boldsymbol{b}}_{2},\xi_{4},C_{4}\delta_{k},V)$-holomorphic.
@@
$\xi:=\min\{\xi_{1},\xi_{2},\xi_{3},\xi_{4},\}\qquad\text{and}\qquad\delta:=\max\Big\{\tilde{C}_{1},C_{2},C_{3},C_{4}K\left\|{f}\right\|_{{\mathcal{K}}_{\varkappa-1}^{r-2}}\Big\}.$
@@
$D\subset{\mathbb{R}}^{2}$ is a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$; $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$ and $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$. For $i=1,2$, the sequences ${\boldsymbol{b}}_{i}:=(b_{i,j})_{j\in{\mathbb{N}}}$ defined as in (5.4) satisfy the condition ${\boldsymbol{b}}_{i}\in\ell^{p_{i}}({\mathbb{N}})$ with $0<p_{1}\leq p_{2}<1$ and $p_{1}<2/3$.
@@
Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\mathbb{R}}^{\infty}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the linear sampling algorithm defined by (5.8), we have that
@@
for an arbitrarily small number $\varepsilon>0$, where the constants $C$ and $C_{\varepsilon}$ is independent of $u$ and $n$.
@@
where the constant $C$ is independent of $u$ and $n$.
@@
Let Assumption 5.1 hold. Let the numbers $\alpha$ and $\beta$ be defined by (5.11). Then for all $n\geq 2$ there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\mathbb{R}}^{\infty}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
for an arbitrarily small number $\varepsilon>0$, where the constants $C$ and $C_{\varepsilon}$ are independent of $u$ and $n$.
@@ 6 Extensions of least squares sampling algorithms
    Let $m:=n$ and $\lceil 20n\log n\rceil$ points be drawn i.i.d. with respect to $\nu$. Let further ${\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{c_{3}n}\in U$ be the subset of points fulfilling [25, Theorem 1] with $c_{3}n\asymp m$ and $\omega_{i}:=\frac{c_{3}n}{\lceil 20n\log n\rceil}(\varrho({\boldsymbol{y}}_{i}))^{-1}$.
@@
This lemma has been formulated in [7]. As mentioned there, the claims (1) and (3) have been proven in [31, Theorem 8] and [25, Theorem 1], respectively. The claim (2) can be proven a similar way based on the result [8, Theorem 6.7].
@@
Let the assumptions of Theorem 2.2 hold. Then for $j=1,2$ and all $n\geq 2$, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in U$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{(j),n}$ defined as in (6.1),
@@
Let the assumptions of Theorem 2.2 hold. Then for $j=1,2$ and all $n\geq 2$, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in U$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{(2),n}$ defined as in (6.1),

### v4

@@ 3 Applications to holomorphic functions
We recall a concept of $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions which has been introduced in [25, Definition 4.1]. For $m\in{\mathbb{N}}$ and a positive sequence ${\boldsymbol{\varrho}}=(\varrho_{j})_{j=1}^{m}$, we put
@@
The following key result on weighted $\ell_{2}$-summability of $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions has been proven in [25, Corollary 4.9].
@@
Assume that $\boldsymbol{b}$} is a non-increasing sequence. Then we have
@@
Let $v$ satisfy Assumption 3.1. Let $q_{i}:=p_{i}/(1-p_{i})$ for $i=1,2$, and $\varepsilon$ be any positive number. Let the number $\beta$ be defined as in (2.25). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.32). Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.29)–(2.31),
@@ 4 Multi-level sparse-grid interpolation algorithms
On the other hand, it holds the following statement which can be proven in a way similar to the proof of [23, Theorem 3.8]. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that
@@ 5 Applications to parametric elliptic PDEs
where $\nu$ is a constant depending on $D$ and $r$. The following result is proven in [25, Theorem 7.8].
@@
From Lemma 5.3 it follows that the operators $P_{k}$, $k\in{\mathbb{N}}_{0}$, are uniformly bounded in $V$. We extend $P_{k}$ as bounded operators in $L_{2}({\mathbb{R}}^{\infty},V;\gamma)$(which with an abuse is denoted again by $P_{k}$) by the formula
@@
$b_{1,j}:=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}},\quad b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\}$ (5.4)
@@
Let $D\subset{\mathbb{R}}^{2}$ be a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$. Let $\alpha=\frac{r-1}{2}$. Let $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$, and $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$. Let $0<p_{1},p_{2}<1$. For $i=1,2$, let the sequences ${\boldsymbol{b}}_{i}:=(b_{i,j})_{j\in{\mathbb{N}}}$ be defined as in (5.4) with ${\boldsymbol{b}}_{i}\in{\color[rgb]{0,0,0}{\ell_{p_{i}}}}({\mathbb{N}})$. Then there exist $\xi>0$ and $\delta>0$ such that for the parametric solution to (5.1) with inputs (1.3)
@@
Proof. The proof of this lemma is a modification of the proof of [25, Proposition 7.12]. We apply [25, Theorem 4.11] on holomorphy of composite functions with $E={\color[rgb]{0,0,0}{L_{\infty}(D)}}$ and $X=V$ to prove the claims (1)–(5).
Step 1. We verify the claim (1). By [25, Proposition 7.12] there exist $\xi_{1}>0$ and $\delta_{1}>0$ such that $u({\boldsymbol{y}})$ is $({\boldsymbol{b}}_{1},\xi_{1},\delta_{1},V)$-holomorphic on the open set
$O_{1}=\{a\in{\color[rgb]{0,0,0}{L_{\infty}(D;{\mathbb{C}})}}\,:\,\rho(a)>0\}\subset{\color[rgb]{0,0,0}{L_{\infty}(D;{\mathbb{C}})}}$
@@
Step 2. Concerning the claim (2), by assumption, $b_{1,j}=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}}$ satisfies ${\boldsymbol{b}}_{1}=(b_{1,j})_{j\in{\mathbb{N}}}\in{\color[rgb]{0,0,0}{\ell_{p_{1}}({\mathbb{N}})\subseteq\ell_{1}({\mathbb{N}})}}$, which corresponds to the assumption (iv) of Theorem [25, Theorem 4.11]. It remains to verify the assumptions (i), (ii) and (iii) of [25, Theorem 4.11] for $P_{k}u$, $k\in{\mathbb{N}}_{0}$.
@@
(iii) For all $a$, $b\in O_{1}$ we have by Lemma 5.2 and [25, (4.21)],
$\left\|{P_{k}{\mathcal{U}}(a)-P_{k}{\mathcal{U}}(b)}\right\|_{V}\leq C\left\|{{\mathcal{U}}(a)-{\mathcal{U}}(b)}\right\|_{V}\leq\left\|{f}\right\|_{H^{-1}(D)}\frac{1}{\min\{\rho(a),\rho(b)\}^{2}}\left\|{a-b}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D;{\mathbb{C}})}}}.$ (5.6)
According to [25, Theorem 4.11] the map
${\mathcal{U}}\mapsto P_{k}{\mathcal{U}}\in{\color[rgb]{0,0,0}{L_{2}}}(U,V;\gamma)$
@@
and verify again the assumptions (i), (ii) and (iii) of [25, Theorem 4.11] with $E=W^{1}_{\infty}(D)\cap{\mathcal{W}}^{s-1}_{\infty}(D)$. First, observe that with
$b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{{\mathcal{W}}^{s-1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{W^{1}_{\infty}(D)}}}\big\},$
@@
${\boldsymbol{b}}_{2}=(b_{2,j})_{j\in{\mathbb{N}}}\in{\color[rgb]{0,0,0}{\ell_{p_{2}}({\mathbb{N}})\hookrightarrow\ell_{1}({\mathbb{N}})}},$
which corresponds to the assumption (iv) of Theorem [25, Theorem 4.11].
@@
$\left\|{({\mathcal{U}}-P_{k}{\mathcal{U}})(a)}\right\|_{V}\leq\delta_{k}\frac{\Big(\left\|{a}\right\|_{W^{1}_{\infty}(D)}+\left\|{a}\right\|_{{\mathcal{W}}^{r-1}_{\infty}(D)}\Big)^{N_{r}+1}}{\rho(a)^{N_{r}+2}},$
[…]
(iii) For every $a$, $b\in O_{2}\subseteq O_{1}$, by (5.6) and [25, (4.21)],
@@
$\leq C\left\|{f}\right\|_{H^{-1}(D)}\frac{2}{\min\{\rho(a),\rho(b)\}^{2}}\left\|{a-b}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D;{\mathbb{C}})}}}.$
We conclude with [25, Theorem 4.11] that there exist $\xi_{4}$ and $C_{4}$ depending on $O_{2}$, $D$ but independent of $k$ such that $u-P_{k}u$ is $({\boldsymbol{b}}_{2},\xi_{4},C_{4}\delta_{k},V)$-holomorphic.
@@
[…]
@@
$D\subset{\mathbb{R}}^{2}$ is a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$; $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$ and $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$. For $i=1,2$, the sequences ${\boldsymbol{b}}_{i}:=(b_{i,j})_{j\in{\mathbb{N}}}$ defined as in (5.4) satisfy the condition ${\boldsymbol{b}}_{i}\in{\color[rgb]{0,0,0}{\ell_{p_{i}}}}({\mathbb{N}})$ with $0<p_{1}\leq p_{2}<1$ and $p_{1}<2/3$.
@@
Let $\varepsilon$ be an arbitrarily positive number. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the linear sampling algorithm defined by (5.8), we have that
@@
where the constants $C$ and $C_{\varepsilon}$ is independent of $n$.
@@
where the constant $C$ is independent of $n$.
@@
Let Assumption 5.1 hold. Let the numbers $\alpha$ and $\beta$ be defined by (5.11). Let $\varepsilon$ be an arbitrarily positive number. Then for all $n\geq 2$ there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
for an arbitrarily small number $\varepsilon>0$, where the constants $C$ and $C_{\varepsilon}$ are independent of $n$ .
@@ 6 Extensions of least squares sampling algorithms
    Let $m:=n$ and $\lceil 20n\log n\rceil$ points be drawn i.i.d. with respect to $\nu$. Let further ${\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{c_{3}n}\in U$ be the subset of points fulfilling [24, Theorem 1] with $c_{3}n\asymp m$ and $\omega_{i}:=\frac{c_{3}n}{\lceil 20n\log n\rceil}(\varrho({\boldsymbol{y}}_{i}))^{-1}$.
@@
This lemma has been formulated in [7]. As mentioned there, the claims (1) and (3) have been proven in [30, Theorem 8] and [24, Theorem 1], respectively. The claim (2) can be proven a similar way based on the result [8, Theorem 6.7].
@@
Let the assumptions of Theorem 2.2 hold. Then for $j=1,2$ and all $n\geq 2$, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{(j),n}$ defined as in (6.1),
@@
Let the assumptions of Theorem 2.2 hold. Then for $j=1,2$ and all $n\geq 2$, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{(2),n}$ defined as in (6.1),